# Diffusion Self-Weighted Guidance for Offline Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning (RL) recovers the optimal policy $\pi$ given historical observations of an agent. In practice, $\pi$ is modeled as a weighted version of the agent’s behavior policy $\mu$, using a weight function $w$ working as a critic of the agent’s behavior. Though recent approaches to offline RL based on diffusion models have exhibited promising results, the computation of the required scores is challenging due to their dependence on the unknown $w$. In this work, we alleviate this issue by constructing a diffusion over both the actions and the weights. With the proposed setting, the required scores are directly obtained from the diffusion model without learning extra networks. Our main conceptual contribution is a novel guidance method, where guidance (which is a function of $w$) comes from the same diffusion model, therefore, our proposal is termed Self-Weighted Guidance (SWG). We show that SWG generates samples from the desired distribution on toy examples and performs on par with state-of-the-art methods on D4RL’s challenging environments, while maintaining a streamlined training pipeline. We further validate SWG through ablation studies on weight formulations and scalability.
@@ 1 Introduction
Offline reinforcement learning (RL) learns a policy from a dataset of an agent’s collected interactions. This setting is commonplace in scenarios such as robotics and healthcare, where real-world exploration with suboptimal agents is costly, dangerous, and thus unfeasible levine2020offlinereinforcementlearningtutorial .
A critical issue when estimating the state-action value function $Q(s,\mathbf{a})$ (Q-values) in offline RL is the unknown effects of actions absent in the dataset, referred to out-of-sample actions. In practice, methods such as iql ; offlinerl_no_ood ; idql refrain from querying out-of-sample actions during training, thus ensuring that the learned policy remains close to the dataset policy, and avoiding distributional shift bear ; bcq ; nair2020awac . These methods typically adopt a weighted behavior policy form given by
@@
Previous approaches learn $\pi(\mathbf{a}|s)$ in a weighted regression procedure weighted_regression ; AWRPeng19 ; nair2020awac ; critic_regularized_regression , using, e.g., a Gaussian model iql ; this approach is severely limited by the expressiveness of the chosen model. For instance, when the dataset contains multi-modal actions, uni-modal policies will certainly fail to capture the agent’s behavior successfully idql ; diff_dice .
Diffusion models (DM) sohl2015deep ; ddpm ; songscore have been successfully applied to sampling from the unnormalized target policy in Eq. (1) janner2022diffuser ; sfbc ; idql ; cep ; diff_dice , while modeling a potentially multi-modal policy dql ; diffusionpolicy . This has been achieved by approximating the behavior policy using a DM, i.e., $\mu_{\theta}(\mathbf{a}|s)\approx\mu(\mathbf{a}|s)$, learning the weight function using external models (see Sec. 2.1 and 2.2), and then constructing a procedure to emulate samples from $\pi(\mathbf{a}|s)$. The main challenge of this approach is the estimation of the intractable score, where the noisy actions and the weights are entangled. This is typically addressed by learning an auxiliary model, referred to as the guidance network cep .
In this work, we propose a fundamentally different approach to diffusion-model-based offline RL. In addition to sampling the actions from the behavior policy $\mu(\mathbf{a}|s)$, we also sample the weights $w(s,\mathbf{a})$ using the same DM. As shown in Sec. 3, this enables a sampling procedure that effectively generates actions from the target policy $\pi(\mathbf{a}|s)$. In contrast to prior approaches that rely on training two distinct models—a DM and an external guidance network—our method removes the need of an external guidance network by embedding the guidance information within the diffusion model itself. We refer to this central contribution as Self-Weighted Guidance (SWG), detailed in Sec.3. Our contributions are summarized as follows
@@
  • A proof-of-concept for SWG’s exact guidance capabilities in toy experiments (Sec. 5.1).
  • Experimental validation for SWG’s competitive performance when compared to the state of the art in D4RL’s challenging environments, with a simpler training setup (Sec. 5.2).
  • Ablation studies exploring SWG’s robustness and scaling properties (Secs. 5.3 and 5.4).
@@ 2.1 Offline reinforcement learning
Consider an agent that, following an unknown behavior policy $\mu(\mathbf{a}|s)$, interacted with the environment generating samples $(s,\mathbf{a},r,s^{\prime})$ recorded in a dataset $\mathcal{D}_{\mu}$. The goal of offline RL is to estimate the optimal policy $\pi(\mathbf{a}|s)$ from $\mathcal{D}_{\mu}$. Modeling the target policy as a weighted behavior policy in Eq. (1) is commonplace across various offline RL methods, some of which are presented next.
Constrained policy optimization peters_2010_constrained_policy ; AWRPeng19 ; nair2020awac regularizes the learned policy to ensure it remains close to the behavior policy by:
@@
Following peters_2010_constrained_policy ; AWRPeng19 ; nair2020awac , it can be shown that the optimal policy $\pi^{*}(\mathbf{a}|s)$ in Eq. (2) has a weighted structure as in Eq. (1), given by:
$\pi^{*}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta Q{(s,\mathbf{a})}},$ (3)
where $Q(s,\mathbf{a})$ represents the Q-values function of the optimal policy. Notice that estimating the target policy Q-values requires querying out-of-sample actions.
Implicit Q-learning (IQL) iql avoids querying out-of-sample actions using an in-sample loss based on expectile regression. This way, it defines an implicit policy that outperforms the behavior policy. Furthermore, IQL employs advantage-weighted regression weighted_regression ; wang_weighted_regression ; AWRPeng19 ; nair2020awac to extract the policy, which can also be interpreted as a weighted behavior policy, given by:
$\pi(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta A{(s,\mathbf{a})}},$ (4)
@@
Implicit Diffusion Q-learning (IDQL) idql generalizes implicit formulations to the use of an arbitrarily convex function $f$:
$V^{*}(s)=\arg\min_{V(s)}\mathbb{E}_{\mathbf{a}\sim\mu(\mathbf{a}|s)}\left[f(Q(s,\mathbf{a})-V(s))\right],$ (5)
@@
$\pi_{\text{imp}}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)w(s,\mathbf{a}),$ (6)
where $w(s,\mathbf{a})=\frac{|f^{\prime}(Q(s,\mathbf{a})-V^{*}(s))|}{|Q(s,\mathbf{a})-V^{*}(s)|}$. Furthermore, IDQL explores different convex functions $f$, such as expectiles iql , quantiles quantilekoenker and exponential functions.
@@ 2.2 Diffusion models in offline RL
Diffusion models (DM) sohl2015deep ; ddpm ; songscore approximate an unknown data distribution [^1] We denote data with the symbol $\mathbf{a}$ (instead of the standard $x$) for consistency, since in our setting the data are the agent’s actions. $q_{0}(\mathbf{a}_{0})$ using a dataset $\mathcal{D}$ by defining a sequence of $k\in\{0,\ldots,K\},K\in\mathbb{N},$ noisy distributions defined as follows:
$q_{k}(\mathbf{a}_{k}|\mathbf{a}_{0})\sim\mathcal{N}(\mathbf{a}_{k};\alpha_{k}\mathbf{a}_{0},\sigma_{k}^{2}\textbf{I}),$ (7)
where $\alpha_{k}$ and $\sigma_{k}$ are predefined noise schedules, s.t., $\mathbf{a}_{K}\sim\mathcal{N}(0,\mathbf{I})$ as $K\to\infty$. Therefore, samples from $q_{0}$ can be obtained by first sampling from $\mathcal{N}(0,\mathbf{I})$, and then solving the reverse diffusion process. In practice, the reverse diffusion is approximated by a model $\epsilon_{\theta}$ that predicts the noise to be removed from a noisy sample $\mathbf{a}_{K}\sim q_{K}$, to produce a target sample $\mathbf{a}_{0}\sim q_{0}$. The training loss for $\epsilon_{\theta}$ is ddpm
$\arg\min_{\theta}\mathbb{E}_{\mathbf{a}_{0},\epsilon_{0},k}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{a}_{k},k)||_{2}^{2},$ (8)
where $\mathbf{a}_{0}\sim q_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, $k\sim\mathcal{U}(1,K)$ and $\mathbf{a}_{k}=\alpha_{k}\mathbf{a}_{0}+\sigma_{k}\epsilon_{0}$. The noise is related to the score function songscore of the intermediate noisy distribution as follows:
$\frac{-\epsilon_{0}}{\sigma_{k}}=\underbrace{\nabla_{\mathbf{a}_{k}}\log{q_{k}(\mathbf{a}_{k}})}_{\text{Score function}}.$ (9)
The weighted behavior policy expression in Eq. (1) admits using DMs in offline RL. To condition the DM on states, the noise prediction network can be made an explicit function of the state, that is, $\epsilon(\mathbf{a}_{k},k)=\epsilon(\mathbf{a}_{k},k,s)$. Previous approaches janner2022diffuser ; sfbc ; idql ; cep ; diff_dice fit a DM $\mu_{\theta}(\mathbf{a}|s)$ to the behavior policy $\mu(\mathbf{a}|s)$, and learn the weights $w(s,\mathbf{a})$ using methods described in Sec. 2.1. Then, samples from the target policy are obtained via resampling or guidance, explained as follows.
Resampling involves first sampling a batch of actions from the learned behavior policy $\mu_{\theta}(\mathbf{a}|s)$, and then resampling them using the weights as probabilities of a categorical distribution idql ; sfbc , in a way inspired by importance sampling. Alternatively, one can adopt a greedy policy that simply selects the action that maximizes the weight, with the aim of improving performance.
Guidance aims to obtain samples from the target policy by modifying the diffusion, adding a score term approximated by an external guidance network $g_{\phi}$ janner2022diffuser ; cep ; diff_dice . The $k$-step score in a diffusion sampling procedure is given by:
@@
where $\{\mathbf{a}_{k}\}_{0}^{K}$ are the noisy versions of the actions, with $\mathbf{a}_{0}$ being the true noiseless action. As shown by cep , an unbiased estimate of the guidance score that effectively produces samples from the target distribution should satisfy:
@@
where the right-hand term is intractable for most policies and weight functions. Guidance methods that are unbiased towards this objective are referred to as exact guidance methods cep .
Diffuser janner2022diffuser , trains the guidance network with a mean-squared-error objective that produces a biased target policy, in the sense that it does not match the form in Eq. (11). Conversely, exact guidance methods such as Contrastive Energy Prediction (CEP) cep use a contrastive learning objective that requires querying a batch of out-of-sample actions for each state in the dataset. Lastly, D-DICE diff_dice , proposes In-sample Guidance Learning (IGL) as an in-sample loss to learn the guidance network and uses a guide-then-select approach to obtain samples.
@@
While exact guidance methods effectively produce samples from the target distribution, they require learning an external network $g_{\phi}$ through an intricate loss function $\mathcal{L}_{g}(\phi)$. This motivates our proposal called Self-Weighted Guidance, which leverages DM’s without resorting to a special training objective to conduct exact guidance.
@@ 3.1 Problem statement
where $\mathbf{a}_{0}\in\mathcal{D}_{A}$ is a data sample, $\mathbf{a}_{1},\ldots,\mathbf{a}_{K}$ are the noisy copies of the data on which the DM is trained, and the subindex $\cdot_{k}$ in the distributions denote the denoising step. Though $\nabla_{\mathbf{a}_{k}}\log q_{k}(\mathbf{a}_{k})$, referred to as intermediate score, is directly provided by the DM, estimating the target score in Eq. (13) is challenging due to the intractability of the right-hand term in the general case cep .
@@ 3.2 Proposal: diffusing over actions and weights
Let us consider a random variable (RV) $Z=(A,W)$ given by the concatenation of $A$, describing the actions in the previous section, and $W$, which is given by passing $A$ through the weight function $w(\cdot)$. Since $w(\cdot)$ is a deterministic function of $A$, the distribution of $Z$ is given by the pushforward measure of the distribution of $A$ through the mapping $T:\mathbf{a}\mapsto(\mathbf{a},w(\mathbf{a}))$. With a slight abuse of notation, we denote the pushforward measures corresponding to $p$ and $q$ in Eq. (12) with the same symbols, i.e., $p(\mathbf{z})=T_{\#}p(\mathbf{a})$ and $q(\mathbf{z})=T_{\#}q(\mathbf{a})$, respectively. Therefore, we have
$p(\mathbf{z})$ $=p(\mathbf{a},w)=p(w|\mathbf{a})p(\mathbf{a})=p(\mathbf{a})\delta_{w=w(\mathbf{a})}$ (14)
$q(\mathbf{z})$ $=q(\mathbf{a},w)=q(w|\mathbf{a})q(\mathbf{a})=q(\mathbf{a})\delta_{w=w(\mathbf{a})},$ (15)
which, combined with Eq. (12), give $p(\mathbf{z})\propto q(\mathbf{z})w$. In this new formulation, $w$ is no longer to be considered as a function of a, but rather as a variable for which there are available observations. For clarity of presentation, we define the following function
@@
The function $\phi_{w}:\mathbb{R}^{d+1}\to\mathbb{R}$ extracts the weight component from the concatenated variable $\mathbf{z}=[\mathbf{a},w]$. That is, $\phi_{w}(\cdot)=\phi_{w}([\mathbf{a},w])=w$.
@@
thus implying that the target distribution for $\mathbf{z}$ follows the same structure as the original distribution in Eq. (12). This motivates jointly learning weights and actions with the same DM. Since the extraction function $\phi_{w}(\mathbf{z})$ is linear, it is feasible to estimate the (otherwise intractable) score within our DM.
We can now implement a DM to sample from the augmented variable $Z$. To this end, we first extend the dataset $\mathcal{D}_{A}=\{(\mathbf{a}^{(i)})\}_{i=1}^{N}$ to $\mathcal{D}_{Z}=\{(\mathbf{a}^{(i)},w^{(i)})\}_{i=1}^{N}\subset\mathbb{R}^{d+1}$, where $w^{(i)}=w(\mathbf{a}^{(i)})$ is a weight observation. Notice that the dataset $\mathcal{D}_{Z}$ provides access to the weight function only on the support of the original data in $\mathcal{D}_{A}$. For consistency, we also denote realizations of $Z$ as $\mathbf{z}^{(i)}=(\mathbf{a}^{(i)},w^{(i)})$.
Therefore, we define a $K$-step diffusion with noise schedule $\{\alpha_{k}\}_{k=1}^{K}$, $\{\sigma_{k}\}_{k=1}^{K}$, and train a noise-prediction DM ${\epsilon}_{\theta}$ for $Z$ using the augmented dataset $\mathcal{D}_{Z}$, minimizing the standard diffusion loss in Eq. (8). In our notation, this is
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{z}_{0},\epsilon_{0},k}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2},$ (17)
where $\mathbf{z}_{0}\sim\mathcal{D}_{Z}$, $\mathbf{z}_{k}=\alpha_{k}\mathbf{z}_{0}+\sigma_{k}\epsilon_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, and $k\sim\mathcal{U}(1,K)$. Algorithm 1 describes our proposed DM training procedure.
@@
     Calculate loss $L(\theta)\approx\frac{1}{B}\sum_{B}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2}$

### v2

@@ Abstract
Offline reinforcement learning (RL) recovers the optimal policy $\pi$ given historical observations of an agent. In practice, $\pi$ is modeled as a weighted version of the agent’s behavior policy $\mu$, using a weight function $w$ working as a critic of the agent’s behavior. Although recent approaches to offline RL based on diffusion models (DM) have exhibited promising results, they require training a separate guidance network to compute the required scores, which is challenging due to their dependence on the unknown $w$. In this work, we construct a diffusion model over both the actions and the weights, to explore a more streamlined DM-based approach to offline RL. With the proposed setting, the required scores are directly obtained from the diffusion model without learning additional networks. Our main conceptual contribution is a novel exact guidance method, where guidance comes from the same diffusion model; therefore, our proposal is termed Self-Weighted Guidance (SWG). Through an experimental proof of concept for SWG, we show that the proposed method i) generates samples from the desired distribution on toy examples, ii) performs competitively against state-of-the-art methods on D4RL when using resampling, and iii) exhibits robustness and scalability via ablation studies.
@@ 1 Introduction
Offline reinforcement learning (RL) learns a policy from a dataset of an agent’s collected interactions. This setting is commonplace in scenarios such as robotics and healthcare, where real-world exploration with suboptimal agents is costly, dangerous, and thus unfeasible (levine2020offlinereinforcementlearningtutorial).
A critical issue when estimating the state-action value function $Q(s,\mathbf{a})$ (Q-values) in offline RL is the unknown effects of the actions absent in the dataset, referred to out-of-sample actions. In practice, methods such as (iql; offlinerl_no_ood; idql) refrain from querying out-of-sample actions during training, thus ensuring that the learned policy remains close to the dataset policy, and avoiding distributional shift (bear; bcq; nair2020awac). These methods typically adopt a weighted behavior policy form given by
@@
Previous approaches learn $\pi(\mathbf{a}|s)$ in a weighted regression procedure (weighted_regression; AWRPeng19; nair2020awac; critic_regularized_regression), using, e.g., a Gaussian model (iql); this approach is severely limited by the expressiveness of the chosen model. For instance, when the dataset contains multimodal actions, unimodal policies will certainly fail to capture the agent’s behavior successfully (idql; diff_dice).
Diffusion models (DM) (sohl2015deep; ddpm; songscore) have been successfully applied to sampling from the unnormalized target policy in equation 1 (janner2022diffuser; sfbc; idql; cep; diff_dice), while modeling a potentially multi-modal policy (dql; diffusionpolicy). This has been achieved by approximating the behavior policy using a DM, i.e., $\mu_{\theta}(\mathbf{a}|s)\approx\mu(\mathbf{a}|s)$, learning the weight function using external models (see Secs. 2.1 and 2.2), and then constructing a procedure to emulate samples from $\pi(\mathbf{a}|s)$. The main challenge of this approach is the estimation of the intractable score, where noisy actions and weights are entangled. This is typically addressed by learning an auxiliary model, referred to as the guidance network (cep).
In this work, we propose a fundamentally different approach to diffusion-model-based offline RL. In addition to sampling actions from the behavior policy $\mu(\mathbf{a}|s)$, we also sample the weights $w(s,\mathbf{a})$ using the same DM. As shown in Sec. 3, this enables a sampling procedure that effectively generates actions from the target policy $\pi(\mathbf{a}|s)$. In contrast to previous approaches that rely on the training of two distinct models —a DM and an external guidance network— our method eliminates the need for an external guidance network by embedding the guidance information within the diffusion model itself. We refer to this central contribution as Self-Weighted Guidance (SWG), detailed in Sec.3.
The main objective of this work is to introduce SWG, a DM over actions and weights for offline RL, and validate its feasibility. Experimentally, we expect a competitive implementation of SWG which performs on par with the state-of-the-art. In particular, our contributions are summarized as follows.
@@
  • A proof of concept for SWG’s exact guidance capabilities in toy experiments (Sec. 6.1).
  • An experimental validation of SWG in D4RL’s challenging environments, where it exhibits competitive performance among recent methods and the ability to leverage a resampling stage (Sec. 6.2).
  • Ablation studies exploring SWG’s robustness and scaling properties (Secs. 6.3, 6.4 and 6.5).
@@ 2.1 Offline reinforcement learning
Consider an agent that, following an unknown behavior policy $\mu(\mathbf{a}|s)$, interacted with the environment generating samples $(s,\mathbf{a},r,s^{\prime})$ recorded in a dataset $\mathcal{D}_{\mu}$. The goal of offline RL is to estimate the optimal policy $\pi(\mathbf{a}|s)$ from $\mathcal{D}_{\mu}$. Modeling the target policy as a weighted behavior policy in equation 1 is commonplace across various offline RL methods, some of which are presented next.
Constrained policy optimization (peters_2010_constrained_policy; AWRPeng19; nair2020awac) regularizes the learned policy to ensure that it remains close to the behavior policy by:
@@
Following (peters_2010_constrained_policy; AWRPeng19; nair2020awac), it can be shown that the optimal policy $\pi^{*}(\mathbf{a}|s)$ in equation 2 has a weighted structure as in equation 1, given by:
$\pi^{*}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta Q{(s,\mathbf{a})}},$
where $Q(s,\mathbf{a})$ represents the Q-values function of the optimal policy. Note that estimating the target policy Q-values requires querying out-of-sample actions.
Implicit Q-learning (IQL) (iql) avoids querying out-of-sample actions using an in-sample loss based on expectile regression. In this way, it defines an implicit policy that outperforms the behavior policy. Furthermore, IQL employs advantage-weighted regression (weighted_regression; wang_weighted_regression; AWRPeng19; nair2020awac) to extract the policy, which can also be interpreted as a weighted behavior policy, given by:
$\pi(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta A{(s,\mathbf{a})}},$
@@
Implicit Diffusion Q-learning (IDQL) (idql) generalizes implicit formulations to the use of an arbitrarily convex function $f$:
$V^{*}(s)=\arg\min_{V(s)}\mathbb{E}_{\mathbf{a}\sim\mu(\mathbf{a}|s)}\left[f(Q(s,\mathbf{a})-V(s))\right],$
@@
$\pi_{\text{imp}}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)w(s,\mathbf{a}),$
where $w(s,\mathbf{a})=\frac{|f^{\prime}(Q(s,\mathbf{a})-V^{*}(s))|}{|Q(s,\mathbf{a})-V^{*}(s)|}$. Furthermore, IDQL explores different convex functions $f$, such as expectiles (iql), quantiles (quantilekoenker) and exponential functions.
@@ 2.2 Diffusion models in offline RL
Diffusion models (DM) approximate an unknown data distribution [^1] We denote data with the symbol $\mathbf{a}$ (instead of the standard $x$) for consistency, since in our setting the data are the agent’s actions. $q_{0}(\mathbf{a}_{0})$ using a dataset $\mathcal{D}$ by defining a sequence of $k\in\{0,\ldots,K\},K\in\mathbb{N},$ noisy distributions defined as follows:
$q_{k}(\mathbf{a}_{k}|\mathbf{a}_{0})\sim\mathcal{N}(\mathbf{a}_{k};\alpha_{k}\mathbf{a}_{0},\sigma_{k}^{2}\textbf{I}),$
where $\alpha_{k}$ and $\sigma_{k}$ are predefined noise schedules, s.t., $\mathbf{a}_{K}\sim\mathcal{N}(0,\mathbf{I})$ as $K\to\infty$. Therefore, samples from $q_{0}$ can be obtained by first sampling from $\mathcal{N}(0,\mathbf{I})$, and then solving the reverse diffusion process. In practice, the reverse diffusion is approximated by a model $\epsilon_{\theta}$ that predicts the noise to be removed from a noisy sample $\mathbf{a}_{K}\sim q_{K}$, to produce a target sample $\mathbf{a}_{0}\sim q_{0}$. The training loss for $\epsilon_{\theta}$ is (ddpm)
[…]
where $\mathbf{a}_{0}\sim q_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, $k\sim\mathcal{U}(1,K)$ and $\mathbf{a}_{k}=\alpha_{k}\mathbf{a}_{0}+\sigma_{k}\epsilon_{0}$. The noise is related to the score function (songscore) of the intermediate noisy distribution as follows:
$\frac{-\epsilon_{0}}{\sigma_{k}}=\underbrace{\nabla_{\mathbf{a}_{k}}\log{q_{k}(\mathbf{a}_{k}})}_{\text{Score function}}.$
Our motivation for employing diffusion models (DMs) in offline RL stems from two main factors: (1) DMs are highly expressive compared to classical generative models (e.g., Mixture Models), allowing them to model multi-modal action distributions (diffusionpolicy); and (2) their score-based formulation enables flexible sampling from conditioned or weighted distributions via guidance. In particular, the weighted behavior in equation 1 can be naturally incorporated using DMs, making them well suited for offline RL.
To condition the DM on the states, the noise prediction network can be made an explicit function of the state, that is, $\epsilon(\mathbf{a}_{k},k)=\epsilon(\mathbf{a}_{k},k,s)$. Previous approaches fit a DM $\mu_{\theta}(\mathbf{a}|s)$ to the behavior policy $\mu(\mathbf{a}|s)$, learn the weights $w(s,\mathbf{a})$ as described in Sec. 2.1, and then sample from the target policy via resampling or guidance; these are explained as follows.
Resampling involves first sampling a batch of actions from the learned behavior policy $\mu_{\theta}(\mathbf{a}|s)$, and then resampling them using weights as probabilities of a categorical distribution (idql; sfbc), in a way inspired by importance sampling. Alternatively, one can adopt a greedy policy that simply selects the action that maximizes the weight with the aim of improving performance.
Guidance aims to obtain samples from the target policy by modifying the diffusion, adding a score term approximated by an external guidance network $g_{\phi}$ (janner2022diffuser; cep; diff_dice). The $k$-step score in a diffusion sampling procedure is given by:
@@
where $\{\mathbf{a}_{k}\}_{0}^{K}$ are the noisy versions of the actions, with $\mathbf{a}_{0}$ being the true noise-free action. As shown by (cep), an unbiased estimate of the guidance score that effectively produces samples from the target distribution should satisfy the following:
@@
where the right-hand term is intractable for most policies and weight functions. Guidance methods that are unbiased towards this objective are referred to as exact guidance methods.
Diffuser (janner2022diffuser), trains the guidance network with a mean-squared-error objective with respect to Q-values, which produces a biased target policy, in the sense that it does not match the form in equation 5. Conversely, exact guidance methods such as Contrastive Energy Prediction (CEP) (cep) use a contrastive learning objective that requires querying a batch of out-of-sample actions for each state in the dataset. Lastly, D-DICE (diff_dice), proposes an in-sample loss to learn the guidance network and uses a guide-then-select approach to obtain samples.
@@
While exact guidance methods effectively produce samples from the target distribution, they require learning an external network $g_{\phi}$ through an intricate loss function $\mathcal{L}_{g}(\phi)$. This motivates our proposal Self-Weighted Guidance, which leverages DM’s without resorting to a special training objective to conduct exact guidance.
@@ 3.1 Problem statement
where $\mathbf{a}_{0}\in\mathcal{D}_{A}$ is a data sample, $\mathbf{a}_{1},\ldots,\mathbf{a}_{K}$ are noisy copies of the data on which the DM is trained, and the subindex $\cdot_{k}$ in the distributions denotes the denoising step. Although $\nabla_{\mathbf{a}_{k}}\log q_{k}(\mathbf{a}_{k})$, referred to as the intermediate score, is provided directly by the DM, estimating the target score in equation 7 is challenging due to the intractability of the right-hand term in the general case (cep).
@@ 3.2 Proposal: diffusing over actions and weights
Let us consider a random variable (RV) $Z=(A,W)$ given by the concatenation of $A$, describing the actions in the previous section, and $W$, which is given by passing $A$ through the weight function $w(\cdot)$. Since $w(\cdot)$ is a deterministic function of $A$, the distribution of $Z$ is given by the push-forward measure of the distribution of $A$ through the mapping $T:\mathbf{a}\mapsto(\mathbf{a},w(\mathbf{a}))$. With a slight abuse of notation, we denote the push-forward measures corresponding to $p$ and $q$ in equation 6 with the same symbols, i.e. $p(\mathbf{z})=T_{\#}p(\mathbf{a})$ and $q(\mathbf{z})=T_{\#}q(\mathbf{a})$, respectively. Therefore, we have
$p(\mathbf{z})$ $=p(\mathbf{a},w)=p(w|\mathbf{a})p(\mathbf{a})=p(\mathbf{a})\delta_{w=w(\mathbf{a})}$
$q(\mathbf{z})$ $=q(\mathbf{a},w)=q(w|\mathbf{a})q(\mathbf{a})=q(\mathbf{a})\delta_{w=w(\mathbf{a})},$
which, combined with equation 6, give $p(\mathbf{z})\propto q(\mathbf{z})w$. In this new formulation, $w$ is no longer to be considered as a function of a, but rather as a variable for which there are available observations. For clarity of presentation, we define the following function
@@
The function $\phi_{w}:\mathbb{R}^{d+1}\to\mathbb{R}$ extracts the weight component from the concatenated variable $\mathbf{z}=[\mathbf{a},w]$. That is, $\phi_{w}:[\mathbf{a},w]\mapsto\phi_{w}([\mathbf{a},w])=w$.
@@
thus implying that the target distribution for $\mathbf{z}$ follows the same structure as the original distribution in equation 6. This motivates jointly learning weights and actions with the same DM. Since the extraction function $\phi_{w}(\mathbf{z})$ is linear, it is feasible to estimate the (otherwise intractable) score within our DM.
We can now implement a DM to sample from the augmented variable $Z$. To this end, we first extend the dataset $\mathcal{D}_{A}=\{(\mathbf{a}^{(i)})\}_{i=1}^{N}$ to $\mathcal{D}_{Z}=\{(\mathbf{a}^{(i)},w^{(i)})\}_{i=1}^{N}\subset\mathbb{R}^{d+1}$, where $w^{(i)}=w(\mathbf{a}^{(i)})$ is a weight observation. Notice that the dataset $\mathcal{D}_{Z}$ provides access to the weight function only on the support of the original data in $\mathcal{D}_{A}$. For consistency, we also denote the realizations of $Z$ as $\mathbf{z}^{(i)}=(\mathbf{a}^{(i)},w^{(i)})$.
Therefore, we define a $K$-step diffusion with noise schedule $\{\alpha_{k}\}_{k=1}^{K}$, $\{\sigma_{k}\}_{k=1}^{K}$, and train a noise-prediction DM ${\epsilon}_{\theta}$ for $Z$ using the augmented dataset $\mathcal{D}_{Z}$, minimizing the standard diffusion loss in equation 3. In our notation, this is
[…]
where $\mathbf{z}_{0}\sim\mathcal{D}_{Z}$, $\mathbf{z}_{k}=\alpha_{k}\mathbf{z}_{0}+\sigma_{k}\epsilon_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, and $k\sim\mathcal{U}(1,K)$. Algorithm 1 describes our proposed DM training procedure. We emphasize that, in practice, Algorithm 1 is implemented with the expectile weight function.
@@
[…]